OpenStudy (anonymous):

MEDAL FOR BEST ANSWER how do you put points into standard form? ex: (1,2) (2,4)

OpenStudy (anonymous):

And can you find standard form in just point? ex: (1,2)

OpenStudy (anonymous):

Given: (-2,-1) → (2, -6). x1 = -2 y1 = -1 x2 = 2 y2 = -6 First, determine slope, m: m = (y2 - y1) / (x2 - x1) m = [-6 - (-1)] / [2 - (-2)] m = (-6 + 1) / (2 + 2) m = -5/4 From the slope-intercept form equation for a line, y = mx + b, we have y = (-5/4)x + b Next, calculate b: b = y / [(-5/4)x] From one point on the line given (-2, -1), substitute -2 for x and -1 for y: b = -1 / [(-5/4)(-2)] b = -1 / (10/4) b = (-1 / )(10/4) b = (-1)(4/10) b = - 4/10 b = -2/5 So, the equation is: y = (-5/4)x - 2/5, or in decimal form: ¯¯¯¯¯¯¯¯¯¯¯¯¯¯ y = -1.25x - 0.4 ¯¯¯¯¯¯¯¯¯¯¯¯¯¯

OpenStudy (anonymous):

I have not given the exact solution of ur question.. Replace ur value of x1,x2,y1 and y2